OpenStudy (anonymous):

Nicholas earned $90.00 for 9 hours of work. How much was earned per hour

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

$10.00 per hour

OpenStudy (anonymous):

how did you get 10.00 ? @heavenfun13

OpenStudy (anonymous):

10*9=90

OpenStudy (anonymous):

but i thought we divide it @heavenfun13

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but wait you just said it was 10 but you siad 10x 9?

OpenStudy (anonymous):

Nicholas earned $90.00 for 9 hours of work 10*9=90 10 per hour

OpenStudy (icecreamyummy):

YES DOLLAR 90.00 / 9 HOURS = 10 dollars PER HOUR